Sol Summary – December 5th

Sol: 8
Summary Title: Two Paths to Hab Ridge
Author’s name: Matt Eby
Mission Status: In-Simulation
Sol Activity Summary: Excursions to Hab Ridge, second interim mirror inspection.
Look Ahead Plan: Drone flight for planning EVA follow-up, soil sample at Somerville Overlook.
Anomalies in work: None
Weather: Calm winds throughout. Cloudy skies early, clearing in the afternoon, with warm temperatures.
Crew Physical Status: Crew is in good shape, nothing of concern to note.
EVA: EVA# 1 led by the Commander, with XO and Health and Safety Officer, traveled by rover to Kissing Camel Ridge, then by foot up Hab Ridge. 2 ½ hour duration. EVA #2 with Crew Engineer and GreenHab Officer returned the mirror sample, then over to North Ridge and ascended Hab Ridge. 2 hour duration on foot.
Reports to be filed: Journalist, Commander, Operations, Photos, EVA report, EVA request, GreenHab
Support Requested: N/A

EVA Report – December 5th

EVA # 10
Author: A. Taylor
Purpose of EVA: Retrieve mirror payload and return it to airlock, test EVA sampling tools (radiation sensor, rake, scoop, sample bags).
Start time: 9:01 AM MT
End time: 11:19 AM MT
Narrative: Ferrone, Braun, and Taylor departed the airlock and headed northeast of the Hab to pick up the mirror coating experiment for inspection and had the payload back in the airlock by 9:15 am. From there the crew took off on the rovers (Taylor and Braun on Curi, and Ferrone on Percy) south on Cow Dung Rd approaching Kissing Camel Ridge. The crew parked the rovers just past the ridge near Robert’s Rock Garden. The trio made their way around the back of the West Kissing Camel Ridge, traversing a large flat area with large boulders on the right. They made their way up red hills and found a path up the side of Hab Ridge to Hab Ridge Rd. The crew took a regolith sample at the top of Hab Ridge. At the halfway mark of the EVA the crew made their way back down the same trail, stopping to take a second regolith sample at the bottom of Kissing Camel Ridge. For both regolith collections the crew used the rake to scrape away the top surface and then scooped the dirt into the handheld sifter, collecting 75 micron size and smaller samples in the special collection bags.
Destination: Stop 1: Retrieve Mirror Payload NE of the hab (not quite to Cow Dung Rd or Marble Ritual) and return to hab. Stop 2: Rovers South on Cow Dung Rd to park near Kissing Camel Ridge – Walk West toward Robert’s Rock Garden and up to Hab Ridge Rd and then back down to Rovers.
Coordinates (use UTM WGS 84): 4249200, 518500
Participants: Commander (Ferrone), XO (Taylor), HSO (Braun)
Road(s) and routes per MDRS Map: Cow Dung Rd. (0110) to Kissing Camel Ridge area
Mode of travel: Rovers (Curi & Percy) and walking

EVA # 11
Author: A. Taylor
Purpose of EVA: Return Mirror Payload and conduct handheld Ham Radio range test.
Start time: 1:53 PM MT
End time: 3:32 PM MT
Narrative: Eby and Kowalski departed the airlock with the mirror surface payload and returned it to its deployment location for one more stint in the Martian environment. From there, they traversed toward the north ridge to test the ham radio operations. The crew traversed up Sagan St toward the North Ridge, and with no obvious path up North Ridge continued westward on Sagan St. toward the intersection with Hab Ridge Rd. Ham radio comms were nominal, even more clear than EVA radios. The crew then made their way back to the deployed mirror location to record the humidity and temperature at the site before returning to the habitat.
Destination: Stop 1: Redeploy Mirror Payload NE of the hab (not quite to Cow Dung Rd or Marble Ritual), Stop 2: Up Sagan St. toward North Ridge stopping before Hab Ridge Rd.
Coordinates (use UTM WGS 84): 4252000, 518500
Participants: Crew Engineer (Kowalski), GreenHab (Eby)
Road(s) and routes per MDRS Map: Hab to Sagan St (1103)
Mode of travel: Rovers Walking

Commander Report – December 5th

Sol: 8
Summary Title: EVAs #10 and #11
Author’s name: Dr. Kristine Ferrone, Commander
Mission Status: Nominal
Commander’s Report:

Crew 269 experienced another day of beautiful weather on Mars with mild temperatures, sunny skies, and light high-altitude clouds. Green Hab Officer made blueberry pancakes for breakfast with bourbon maple syrup. After a morning planning meeting, CDR, XO, and HSO conducted EVA #10. The first objective of the EVA was to traverse on foot to retrieve the mirror coating experiment samples from their exposure location across Cow Dung Rd. from Marble Ritual and return the samples to the airlock for analysis under microscope in the Science Dome by Engineer. The EVA then proceeded on rovers to Kissing Camel Ridge, then on foot up to Hab Ridge Rd. for the purpose of regolith sample collection and radiation environment monitoring. After EVA #10 crew returned to the Hab, Engineer and Green Hab Officer embarked on EVA #11, traversing on foot to return the mirror coating experiment samples to their exposure location across Cow Dung Rd. from Marble Ritual. The EVA then proceeded on foot to Sagan St. for the purpose of a handheld ham radio range test with the Hab. The handheld ham radios maintained clear communications on the 2 meter band at a further distance from the Hab than was possible using the EVA radios. During the afternoon EVA, HSO conducted in-depth troubleshooting of the Hab toilet (details can be found in the operations report for today), and Technology Officer conducted troubleshooting of the VR/AR headset in preparation for future EVAs. In the early evening, Green Hab Officer harvested microgreens, HSO collected crew health stats, and the crew took turns completing their exercise protocols. For dinner, the crew enjoyed chicken jambalaya with homemade sourdough bread and a microgreen salad. After dinner, the crew met to discuss prioritizing the remaining mission objectives to ensure maximum science return with the time remaining.

Operations Report – December 5th

SOL: 8

Name of person filing report: Ashley Kowalski

Non-nominal systems: Toilet Issues

Notes on non-nominal systems:

The Health and Safety Officer, Barbara Braun, performed troubleshooting of the toilet. The holding tank for the toilet has been filling more quickly than expected and smelling bad for longer than expected after emptying. Upon advice of Tier 1 support, the HSO first emptied the holding tank into the septic tank, then flushed the holding tank with approximately two gallons of water. Then the HSO examined the inside of the tank with a headlamp and used a stick to estimate the level of material remaining in the tank after flushing. There were approximately two inches of watery sludge in the bottom of the holding tank even with the valve to the septic tank open. Upon the advice of Tier 1 support, the HSO then closed the septic valve, filled the tank full of fresh water until the water came up to the level of the bowl, and emptied the holding tank again by opening the septic valve. The results were essentially the same; approximately 1-2 inches of watery sludge still remained in the bottom of the holding tank even after the full water flush. Tier 1 support then advised the HSO to fill the tank with approximately five gallons of fresh water and add triple the usual amount of digestive. HSO accomplished this. Note that this activity used up one of the remaining jugs of digestive; we still have a full jug available. Tier 1 support will investigate further and consider servicing the holding tank in a break between rotations.


Spirit rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Opportunity rover used: No
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Curiosity rover used: Yes
Hours: 215.3 (Start), 215.6 (End)
Beginning charge: 100%
Ending charge: 70%
Currently charging: No.

Perseverance rover used: Yes
Hours: 250.5 (Start), 250.7 (End)
Beginning charge: 100%
Ending charge: 56%
Currently charging: No.

General notes on rovers: No additional notes/comments.

Summary of Hab operations:
WATER USE: 300 used (~250 gallons remaining – This is my best guess, but I am requesting Mission Support to let us know a more accurate value, if possible.)
Water (static tank): The anomalous toilet situation today caused the crew to use additional water resources. Tier 1 support provided an additional 120 gallons to make up for this loss.
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes. Special instructions that were provided to the crew by Mission Support for the anomalous toilet situation were followed. Please reference the “Non-nominal systems” section for details.

Summary of internet: No issues with the internet. Connection is good.

Summary of suits and radios:

The Crew Engineer performed the additional steps that were suggested to her by Mr. Scott Davis from the EVA suit support team. The Crew Engineer fully charged suits 3, 4, and 6 so that all battery charges were >12 V. She then ran each of those suits for about an hour. Suits 3 and 4 maintained a voltage >12 V. However, Suit 6 dropped to about 10.41 V. She left Suit 6 recharging for several hours, yet the battery did not charge up again to >12 V. In fact, after several hours of charging, the suit only charged to ~10.47 V. The suit seemed to still be running normally at that voltage. Thus, the suit seems to still be functional, but there might indeed be a slight battery issue with Suit 6. Please advise.

Finally, the two walkie talkies mentioned previous reports (one from the upstairs set and one from the downstairs set) are still not working.

Summary of GreenHab operations: Additional information can be found in the GreenHab Report.
WATER USE: ~39 gallons (~5 gallons today)
Heater: On
Supplemental light: On
Harvest: 17 grams of microgreens and 11 grams of volunteer plants

Summary of ScienceDome operations: Microscope analysis of the mirror samples were performed using The Aerospace Corporation microscope in the ScienceDome today.
Dual split: The heat is on in the ScienceDome and set to 65 F. Please note that as the Crew Engineer was performing mirror sample inspections today in the ScienceDome, it was quite cold still.

Summary of RAM operations: The multimeter from the RAM was used to check the voltages of the two-piece EVA suits during today’s inspections. Additionally, a large wooden paint stick (to assist with the toilet issues), the pulse oximeter, rubbing alcohol, and several sharpies were retrieved from the RAM.

Summary of any observatory issues: No observatory issues.

Summary of health and safety issues: No health and safety issues.

Questions, concerns and requests to Mission Support: Heat furnace filter has not yet been installed. Please advise if this is something that the Crew Engineer needs to do.

Journalist Report – December 5th

Sol: 8
Author: Barbara Braun, Health and Safety Officer
Title: Kissing Camels and Ham Radios

Sol 8 dawned slightly overcast on Mars, but the fresh pancakes made by Green Hab Officer Matt Eby made up for any lack of sunshine. Fortified by this excellent breakfast, Commander Kristine Ferrone, XO Alli Taylor, and Health and Safety Officer Barbara Braun left on the first EVA of the day. The team retrieved the mirror coating experiment for analysis by Crew Engineer Ashley Kowalski, and after stowing the payload in the airlock, headed toward Kissing Camel Ridge on a pair of crew rovers. South of Kissing Camel Ridge they parked Curiosity and Perseverance and ventured out on foot to the west of the road. After crossing a flat plain strewn with large rocks, they ascended through several foothills to Hab Ridge Road, where they were rewarded with beautiful views and multiple samples of regolith. After admiring the views and examining a large cache of small shell-like rocks, the team descended back to the rovers and returned to the Hab.

Meanwhile, Ashley completed her analysis of the mirror samples and prepared for an EVA of her own with Green Hab Officer Matt Eby. Ashley and Matt returned the mirror sample payload to its location for one final round of exposure to the elements. The duo then proceeded up Sagan Street to conduct tests of ham radio communications with Kristine and Barbara back at the Hab. Space exploration may seem glamorous, but there is always the element of the mundane: HSO Barbara Braun spent a good portion of the afternoon troubleshooting the Hab toilet. After this unpleasant task was successfully concluded, and afternoon exercise sessions were completed, the crew settled in for the much more enjoyable activity of eating a dinner of jambalaya, sourdough, and fresh greens harvested by Matt from the Green Hab.

GreenHab Report – December 5th

GreenHab Officer: Matthew Eby
Environmental control: Heater on, door closed, 20% humidity.
Average temperatures: 72-deg F at time of watering. Range prior 24 hours: 61-deg F – 81-deg F
Hours of supplemental light: 4 hours
Daily water usage for crops: Approximately 5 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank: Approximately 205 gallons (Based on a 250 gallon capacity, visually 80-85% full)
Time(s) of watering for crops: 4:30 PM
Changes to crops: None
Narrative: Today the positions of the pea plants and the tomatoes were swapped in an effort to move the tomato plants closer to the heat and the pea plants further away from the heat.
Harvest: 17g microgreens, 11g of “volunteer” plants
Support/supplies needed: None needed at this time.

Copyright © The Mars Society. All rights reserved. | Main Site